OpenStudy (anonymous):

Write an equation of a line that is parallel to the given line and passes through the given point. y = -3x +2 ; (2,-6) Please show steps in to getting the answer. Thankyou.

jimthompson5910 (jim_thompson5910):

y = mx+b y = -3x+b -6 = -3(2)+b Now solve for b

OpenStudy (anonymous):

what do you get after the b? because b becomes 0

jimthompson5910 (jim_thompson5910):

-6 = -3(2)+b -6 = -6+b -6+6 = b 0 = b So b = 0 which means y = mx+b becomes y = -3x+0 y = -3x

OpenStudy (anonymous):

Thankyou

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (rayford):

wow..